# Photometric and spectroscopic variability of blue supergiant $\rho$ Leo

### v1

@@ 3.2.1 Analysis of the K2 light curve
As mentioned in Sect. 2.2, the light curve for K2 was digitized from 10.1093/mnras/sty308. However, we have applied different frequency analysis methods. The results are presented in Table 3. First of all, we implemented a GLS for K2 photometry. The length of K2 photometry was 78.64 days, hence we can take a range of periods for analysis up to 39.32 days. The power spectrum is presented in Fig. 2. We note that the vast majority of frequencies lie in the long-wave region (periods over 1.25 days).
The clear predominance of low frequencies visible in the Lomb-Scargle spectrum is of physical origin and correlates with the stochastic low-frequency (SLF) variability intrinsic to massive stars (2019NatAs...3..760B; 2020A&A...640A..36B; 2023Ap&SS.368..107B). SLF variability in the photometry of massive stars is currently believed to be inherent to the stars themselves, arising from internal gravitational waves and turbulent convection.
Table 3: Periods detected in the K2 light curve. [^8] Results of the analysis of three methods: GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power. Asterisk (*) indicates an uncertainty less than $0.04$ days. In this and the following tables, persistent periods detected in all datasets are highlighted in bold.
@@
Figure 2: Lomb-Scargle power spectrum of the K2 photometry. The first six most powerful frequencies are shown. The dominant frequency (0.0405 d-1) refers to period of 24.7 days.
Figure 3: Phase diagram of K2 photometry with the period of 24.85 days.
Figure 4: WWZ scalogram for K2 photometry. Wavelet power is indicated by color. The lower bright stripe corresponds to the period of $\sim$25 days. The black solid line on the right panel shows the time-averaged wavelet power.
To verify the periods obtained and find the exact value of the periods, the GLSp method was applied to the data. In Table 3, we present a set of frequencies and sub-harmonics that we found. We can notice that frequency $F_{5}\sim F_{2}+F_{4}$ and frequency $F_{6}\sim F_{4}+F_{5}$. Figure 3 shows phase diagram for the most significant period, 24.85 $\pm$ 0.61 days, which is consistent within uncertainties with the 26.8-day period reported by 10.1093/mnras/sty308. The second strongest peak corresponds to period 16.07 ± 0.25 days, see discussion in Sec. 4.
In addition, we performed WWZ analysis and obtained a scalogram, Fig. 4, to study the change in frequency and amplitude over time. The 25.3-day period is clearly visible as the brightest stripe on the bottom of the WWZ scalogram. This period is not stable; it increases towards the end of the observations. Notice also that the power of this period is more pronounced at the beginning. The 10-day period also exists, but it is much weaker. The shorter periods appear as spots.
@@ 3.2.2 Analysis of the TESS light curve
The first results of our analysis are shown in the GLS frequency scalogram presented in Fig. 5. This figure offers a detailed view of the dominant frequencies present in the light curve and also includes the FAP level. We identify the set of significant frequencies, up to 0.63 d-1, listed in Table 4. Note that the evident SLF pattern is similar to the K2 photometry (Fig. 2). The dominant frequency is around 0.055 d-1, or the period of $P\sim 18.2$ days. However, not all significant frequencies are independent. Many signals are sub-harmonics or combinations of the independent frequencies: we can notice that $F_{4}\sim F_{1}+F_{3}$ and $F_{5}\sim 4\times F_{1}$.
Table 4: Periods detected in TESS photometry. [^9] Results of the analysis using three different methods: GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power.
@@
Figure 5: Lomb-Scargle power spectrum of the TESS photometry. The first seven most powerful frequencies are shown. The dominant frequency of 0.055 d-1 refers to period of 18.2 days.
@@
Figure 6: One of the most powerful periods from GLS prewhitened analysis for the TESS sector 46. Prewhitened data phased to the indicated period.
@@
Figure 7: WWZ scalogram for TESS photometry. Wavelet power is indicated by colour. The lower bright stripe corresponds to the period of $\sim$18 days. Next to it above, appears faint stripe, which corresponds to period of $\sim$11 days. Shorter periods – 8, 5.4 days appear as non-permanent spot-like regions in upper part on the colour scalogram. The black solid line on the right panel shows the time-averaged wavelet power.
@@
By constructing a phase diagram, we can illustrate how the period aligns with our data. Moreover, it enables us to pinpoint a more accurate value for the period. Hence, the next step is GLSp. In Table 4 we present a set of frequencies that we found. The dominant frequency is approximately 0.054 d-1, corresponding to a period of $P\sim 18.5$ days. The frequency $F_{4}\sim 0.092$ d-1 ($P_{4}\sim 11$ days) is close to the GLS-identified $F_{2}$. The analysis was carried out using data from sectors 45 and 46 combined. When analysed separately, sector 46 data give a slightly longer period of $12.1\pm 1.3$ days, consistent within uncertainties. For clarity of presentation, we show the phase diagram for sector 46 using the period specific to that sector (Fig. 6), where the signal appears more clearly. The remaining signals can be interpreted as sub-harmonics or combinations of independent frequencies, e.g. $F_{3}\sim 2\times F_{1}$ and $F_{5}\sim F_{2}+F_{3}$.
A weak signal at the frequency corresponding to the period of $26\pm 3$ days, obtained using the GLSp method, is close to the dominant frequency reported by 10.1093/mnras/sty308 based on K2 photometry, and also identified in our analysis of the same dataset.
@@
Finally, we present the WWZ scalogram to understand the temporal evolution of the frequencies and find quasi-periodic events (Figure 7). We found that the most significant frequencies are in the low frequency range. Therefore, we provide here the resulting diagram with a range from 0.038 to 0.25, which corresponds to the period range from 4 to 26.22 days. The upper limit of 26.2 days is linked to the time series length of TESS photometry (sectors 45 and 46 in total had an observation period of 52.44 days).
@@
The strongest signal, with a period of 18.6 days, is relatively stable: its power slowly decreases toward the end of the observing sequence, while the frequency shows a slight increase. Although less sharply defined, the 11.4-day period remains consistently present throughout. The 7.8-day period is clearly detected at the beginning of the observations, but its power declines rapidly, becoming almost undetectable by the end of the sequence.
Additional shorter periods appear intermittently, mainly at the beginning or end of the observing sequence.
@@ 3.2.3 Analysis of the spectroscopic data
The spectroscopic frequency analysis is based on the He I  6678.151 Å line (from here on referred to as He I line). We examined the temporal variations in the profile of this line and performed an analysis of its first moment $\langle v_{\rm 1}\rangle$ (the so-called centroid), which represents changes in the radial velocity of a star, in accordance with the Equation 5.67 of 2010aste.book.....A. To investigate the frequency content of the spectral line variability, we adopted an average radial velocity of the system’s centre of mass, $v_{\rm rad}$ = 42 km s-1, as the zero point on the wavelength scale. The same methods as for the light curve analyses, GLS, GLSp, and WWZ, were exploited.
@@ 3.2.3.A Season 2017
In Fig. 17, we present the first moment measurements for Season 2017 from our 369 spectra collected during 20 nights in period 4 January 2017 to 16 May 2017.
In that Season, we did not have large gaps between observational runs in the second half of the observing period. Therefore, we can take quite a long frequency range for analysis. The lower limit of 0.0076 d-1 corresponds to $1/T$, where T is the observing period (131.8 days). The upper frequency limit was taken as 1.2 d-1 because no high frequencies with a sufficient signal-to-noise ratio were found.
First, the GLS method was applied. The power spectrum is presented in Fig. 8 and the corresponding periods in Table 5. One can notice that the general view of the periodogram differs from the same analysis of photometric data. In addition, we can notice several periods close to 1 day. Most probably, these periods are connected with observational selection because we had a number of observing nights in a row.
Table 5: Periods detected in the Season 2017.101010Results of the analysis using three different methods: GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power.
@@
Figure 8: GLS power spectrum for the Season 2017.
@@
Figure 9: Phase diagrams of the Season 2017 obtained with the GLSp method.
@@
Figure 10: Result of the WWZ analysis for Season 2017. The black solid line on the right panel shows the time-averaged wavelet power.
@@
Secondly, to derive periods and their powers, we performed the GLSp. The periods found are listed in Table 5. We can notice that the fourth frequency $F_{\mathrm{4}}$ (0.1667) is roughly equal to the sum of $F_{\mathrm{1}}$ (0.0582) and $F_{\mathrm{3}}$ (0.0911), also $F_{\mathrm{5}}$ (0.3157) $\sim$ the twofold frequency $F_{\mathrm{4}}$. Phase folded plots of the two persistent periods ($\sim$ 17 and 11 days) are shown in Fig. 9. Note that the folded velocity curve with a period of 17.21 days is not sinusoidal but has a sawtooth-like shape, whereas the curve with a period of 10.98 days can be well approximated by a sine wave.
@@
It is difficult to determine the best-fit curve by eye. To quantify the asymmetry of the phase curve, we compared two simple models: a pure sinusoid and a piecewise linear (“saw-tooth”) function with the same period. Both models were fitted to the phase-folded velocity data using least-squares minimization. The goodness of each fit was evaluated using the root-mean-square (RMS) of the residuals, as well as the Akaike Information Criterion (AIC) 1974ITAC...19..716A and the Bayesian Information Criterion (BIC) 10.1214/aos/1176344136, consistent with a small but measurable asymmetry between the rising and declining branches. Differences for the two fits for the phase curve with a period of 17.21 days are: $\Delta\text{AIC}\sim 48$ and $\Delta\text{RMS}\sim 0.16$, which means that saw-tooth fitting better describes behaviour of the phase curve.
@@
Lastly, we performed the WWZ analysis to understand the behaviour of frequencies and find quasi-periodic events (Fig.10). The range of periods was taken from 4 to 61.4 days, because this range appeared to be the most informative for this type of analysis. In addition to several long periods in the lower part of the plot, one short period is also visible in the form of a fuzzy spot in the upper left part. The results obtained from the time-averaged power spectrum are given in Table 5. As this method gives only rough frequency estimates, the uncertainties are much higher than for the GLSp method. It is clearly visible on the scalogram that two periods, $\sim 6$ days and $\sim 7$ days, presented as a yellow spot on the right side near 1870 HJD, do not exist all the time, but only at the end of the Season. Several long periods are at the bottom, and the strongest of them is 17.4 days. The period of $\sim 11.7$ days, already identified in other analyses, is also present in this dataset. We note that while most of the results obtained by the WWZ analysis agree well with the GLS results (see Table 5), the two longest periods, 22 and $\sim 29$ days, are not detected by any of the GLS analyses. These two periods on the scalogram have an unclear shape and are not constant in time. This suggests that they are likely false. Consequently, we do not consider them further in our analysis.
@@ 3.2.3.B Season 2022
Spectroscopic observations in Season 2022 were made mostly right after the TESS photometry. However, one observing night was just before TESS, 1 November 2021, and two during TESS observations, 21 November 2021 and 2021-12-09. In Fig. 18, we present the first moment measurements for that Season. As shown in the figure, we obtained a large number of spectra during the Season 2022 – 1106 over 39 nights – which enabled us to perform frequency analysis over a broad range with high confidence. The lower limit 0.005 d-1 corresponds to $1/T$, where T is the observing period (194.7 days), and the upper limit is 1.2 d-1.
As with other analyses, we started with the simple GLS periodogram without prewhitening. The corresponding power spectrum is presented in Fig. 11(a). We determined that all significant frequencies lie below 1.2 d-1. As in the previous analysis, we determine the first seven frequencies, see Table 6. The period of $\sim 10$ days is also observed taking into account the uncertainties in this Season. The most powerful period of the spectroscopic data is $\sim 16.2$ days. Notice that a general view of the periodogram is different from the power spectrum of the TESS photometry (Fig. 5).
Table 6: Periods detected in the Season 2022.111111Results of the analysis using three different methods: GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power.
@@
To derive periods and build phase-folded diagrams, we perform the GLSp, with the resulting periods listed in Table 6. We can notice that $F_{\mathrm{5}}$ $\sim$ $F_{\mathrm{2}}$ + $F_{\mathrm{3}}$. In Fig. 11(b), we present a phase plot with the first most powerful period we have found. Note a sawtooth-like shape of the curve.
@@
Figure 12: Result of the WWZ analysis for Season 2022. Wavelet power is indicated by color. The black solid line on the right panel shows the time-averaged wavelet power.
Lastly, we obtained a very interesting result using WWZ analysis (Fig. 12 and Table 6). The frequency range was taken from 0.01 to 0.25 d-1. The lowest and strongest frequency on the scalogram corresponds to the period $\sim 16.4$ days. According to the scalogram, this period was very pronounced and existed throughout the entire observation Season and slightly increases by the end of the Season. The period $\sim 11$ days also exists in the scalogram of the Season 2022, and here the value is equal to 10 $\pm$ 1 days, although this period does not exist all the time and increases toward the end, according to the scalogram.

### v2

@@ 3.2.1 Analysis of the K2 light curve
As mentioned in Sect. 2.2, the light curve for K2 was digitized from 10.1093/mnras/sty308. However, we applied different frequency analysis methods. The results are presented in Table 3. First, we implemented a GLS for the K2 photometry. The length of the K2 photometry was 78.64 days, and hence we can take a range of periods for analysis of up to 39.32 days. The power spectrum is presented in Fig. 2. We note that the vast majority of frequencies lie in the long-wave region (periods over 1.25 days).
The clear predominance of low frequencies visible in the Lomb–Scargle spectrum is of physical origin and correlates with the stochastic low-frequency (SLF) variability intrinsic to massive stars (2019NatAs...3..760B; 2020A&A...640A..36B; 2023Ap&SS.368..107B). SLF variability in the photometry of massive stars is currently believed to be inherent to the stars themselves, arising from internal gravitational waves and turbulent convection.
Table 3: Periods detected in the K2 light curve. [^8] Results of the analysis of three methods: GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power. An asterisk (*) indicates an uncertainty of less than $0.04$ days. In this and the following tables, persistent periods detected in all datasets are highlighted in bold.
@@
Figure 2: Lomb–Scargle power spectrum of the K2 photometry. The first six most powerful frequencies are shown. The dominant frequency (0.0405 d-1) refers to a period of 24.7 days.
Figure 3: Phase diagram of K2 photometry with a period of 24.85 days.
Figure 4: WWZ scalogram for K2 photometry. The wavelet power is indicated by the colour (see colour bar at left). The lower bright stripe corresponds to the period of $\sim$25 days. The black solid line in the right panel shows the time-averaged wavelet power.
To verify the periods obtained and find the exact value of the periods, the GLSp method was applied to the data. In Table 3 we present a set of frequencies and the sub-harmonics that we found. We note that frequency $F_{5}\sim F_{2}+F_{4}$ and frequency $F_{6}\sim F_{4}+F_{5}$. Figure 3 shows the phase diagram for the most significant period, 24.85 $\pm$ 0.61 days, which is consistent within the uncertainties with the 26.8-day period reported by 10.1093/mnras/sty308. The second-strongest peak corresponds to a period of 16.07 ± 0.25 days (see discussion in Sect. 4).
In addition, we performed a WWZ analysis and obtained a scalogram (Fig. 4) to study the change in frequency and amplitude over time. The 25.3-day period is clearly visible as the brightest stripe at the bottom of the WWZ scalogram. This period is not stable; it increases towards the end of the observations. We also note that the power of this period is more pronounced at the beginning. The 10-day period also exists, but it is much weaker. The shorter periods appear as spots.
@@ 3.2.2 Analysis of the TESS light curve
The first results of our analysis are shown in the GLS frequency scalogram presented in Fig. 5. This figure offers a detailed view of the dominant frequencies present in the light curve and also includes the FAP level. We identify the set of significant frequencies, up to 0.63 d-1; they are listed in Table 4. We note that the evident SLF pattern is similar to the K2 photometry (Fig. 2). The dominant frequency is around 0.055 d-1, or the period of $P\sim 18.2$ days. However, not all significant frequencies are independent. Many signals are sub-harmonics or combinations of the independent frequencies; we note that $F_{4}\sim F_{1}+F_{3}$ and $F_{5}\sim 4\times F_{1}$.
Table 4: Periods detected in TESS photometry. [^9] Results of the analysis using three different methods, GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power.
@@
Figure 5: Lomb–Scargle power spectrum of the TESS photometry. The first seven most powerful frequencies are shown. The dominant frequency of 0.055 d-1 refers to a period of 18.2 days.
@@
Figure 6: One of the most powerful periods from GLS pre-whitened analysis for the TESS sector 46. Pre-whitened data phased to the indicated period.
@@
Figure 7: WWZ scalogram for TESS photometry. The wavelet power is indicated by the colour (see colour bar at left). The lower bright stripe corresponds to the period of $\sim$18 days. A faint stripe appears above it, which corresponds to a period of $\sim$11 days. Shorter periods of 8 and 5.4 days appear as non-permanent spot-like regions in the upper part on the colour scalogram. The black solid line in the right panel shows the time-averaged wavelet power.
@@
By constructing a phase diagram, we can illustrate how the period aligns with our data. Moreover, it enables us to pinpoint a more accurate value for the period. Hence, the next step is GLSp. In Table 4 we present the set of frequencies that we found. The dominant frequency is approximately 0.054 d-1, corresponding to a period of $P\sim 18.5$ days. The frequency $F_{4}\sim 0.092$ d-1 ($P_{4}\sim 11$ days) is close to the GLS-identified $F_{2}$. The analysis was carried out using data from sectors 45 and 46 combined. When analysed separately, sector 46 data give a slightly longer period of $12.1\pm 1.3$ days, consistent within the uncertainties. For clarity of presentation, we show the phase diagram for sector 46 using the period specific to that sector (Fig. 6), where the signal appears more clearly. The remaining signals can be interpreted as sub-harmonics or combinations of independent frequencies (e.g. $F_{3}\sim 2\times F_{1}$ and $F_{5}\sim F_{2}+F_{3}$). A weak signal at the frequency corresponding to the period of $26\pm 3$ days, obtained using the GLSp method, is close to the dominant frequency reported by 10.1093/mnras/sty308 based on K2 photometry, and also identified in our analysis of the same dataset.
@@
Finally, we present the WWZ scalogram to understand the temporal evolution of the frequencies and find quasi-periodic events (Figure 7). We found that the most significant frequencies are in the low-frequency range. Therefore, we provide here the resulting diagram with a range from 0.038 to 0.25, which corresponds to the period range from 4 to 26.22 days. The upper limit of 26.2 days is linked to the time series length of TESS photometry (sectors 45 and 46 in total had an observation period of 52.44 days).
@@
The strongest signal, with a period of 18.6 days, is relatively stable. Its power slowly decreases towards the end of the observing sequence, while the frequency shows a slight increase. Although less sharply defined, the 11.4-day period remains consistently present throughout. The 7.8-day period is clearly detected at the beginning of the observations, but its power declines rapidly, becoming almost undetectable by the end of the sequence. Additional shorter periods appear intermittently, mainly at the beginning or end of the observing sequence.
@@ 3.2.3 Analysis of the spectroscopic data
The spectroscopic frequency analysis is based on the He I  6678.151 Å line (hereafter the He I line). We examined the temporal variations in the profile of this line and performed an analysis of its first moment $\langle v_{\rm 1}\rangle$ (called the centroid), which represents changes in the radial velocity of a star, in accordance with Eq. 5.67 in 2010aste.book.....A. To investigate the frequency content of the spectral line variability, we adopted an average radial velocity of the system’s centre of mass, $v_{\rm rad}$ = 42 km s-1, as the zero point on the wavelength scale. The same methods as for the light curve analyses, GLS, GLSp, and WWZ, were exploited.
@@ 3.2.3.A Season 2017
In Fig. 17, we present the first-moment measurements for Season 2017 from our 369 spectra collected during 20 nights in the period 4 January 2017 to 16 May 2017. In that season we did not have large gaps between observational runs in the second half of the observing period. Therefore, we can take quite a long frequency range for analysis. The lower limit of 0.0076 d-1 corresponds to $1/T$, where T is the observing period (131.8 days). The upper frequency limit was taken as 1.2 d-1 because no high frequencies with a sufficient signal-to-noise ratio were found.
First, the GLS method was applied. The power spectrum is presented in Fig. 8 and the corresponding periods in Table 5. We note that the general view of the periodogram differs from the same analysis of photometric data. In addition, several periods close to 1 day can be seen. Most probably, these periods are connected with observational selection because we had a number of observing nights in a row.
Table 5: Periods detected in Season 2017.101010Results of the analysis using three different methods, GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power.
@@
Figure 8: GLS power spectrum for Season 2017.
@@
Figure 9: Phase diagrams of Season 2017 obtained with the GLSp method.
@@
Figure 10: Result of the WWZ analysis for Season 2017. The black solid line in the right panel shows the time-averaged wavelet power.
@@
Secondly, to derive the periods and their powers, we performed the GLSp. The periods found are listed in Table 5. We note that the fourth frequency $F_{\mathrm{4}}$ (0.1667) is roughly equal to the sum of $F_{\mathrm{1}}$ (0.0582) and $F_{\mathrm{3}}$ (0.0911), also $F_{\mathrm{5}}$ (0.3157) $\sim$ the twofold frequency $F_{\mathrm{4}}$. Phase-folded plots of the two persistent periods ($\sim$ 17 and 11 days) are shown in Fig. 9. We note that the folded velocity curve with a period of 17.21 days is not sinusoidal but has a sawtooth-like shape, whereas the curve with a period of 10.98 days can be well approximated by a sine wave.
@@
It is difficult to determine the best-fit curve by eye. To quantify the asymmetry of the phase curve, we compared two simple models: a pure sinusoid and a piecewise linear (sawtooth) function with the same period. Both models were fitted to the phase-folded velocity data using least-squares minimization. The goodness of each fit was evaluated using the root mean square (RMS) of the residuals, as well as the Akaike information criterion (AIC) 1974ITAC...19..716A and the Bayesian information criterion (BIC) 10.1214/aos/1176344136, consistent with a small but measurable asymmetry between the rising and declining branches. The differences for the two fits for the phase curve with a period of 17.21 days are $\Delta\text{AIC}\sim 48$ and $\Delta\text{RMS}\sim 0.16$, which means that sawtooth fitting better describes the behaviour of the phase curve.
@@
Lastly, we performed the WWZ analysis to understand the behaviour of the frequencies and find the quasi-periodic events (Fig.10). The range of periods was taken from 4 to 61.4 days because this range appeared to be the most informative for this type of analysis. In addition to several long periods in the lower part of the plot, one short period is also visible in the form of a fuzzy spot in the upper left part. The results obtained from the time-averaged power spectrum are given in Table 5. As this method gives only rough frequency estimates, the uncertainties are much higher than for the GLSp method. It is clearly visible on the scalogram that two periods, $\sim 6$ days and $\sim 7$ days that presented as a yellow spot on the right side near 1870 HJD, do not exist all the time, but only at the end of the season. Several long periods are at the bottom, and the strongest of them is 17.4 days. The period of $\sim 11.7$ days, already identified in other analyses, is also present in this dataset. We note that while most of the results obtained by the WWZ analysis agree well with the GLS results (see Table 5), the two longest periods, 22 and $\sim 29$ days, are not detected by any of the GLS analyses. These two periods on the scalogram have an unclear shape and are not constant in time. This suggests that they are likely false. Consequently, we do not consider them further in our analysis.
@@ 3.2.3.B Season 2022
Spectroscopic observations in Season 2022 were made mostly right after the TESS photometry. However, one observing night was just before TESS, 1 November 2021, and two during the TESS observations, 21 November 2021 and 09 December 2021. In Fig. 18, we present the first-moment measurements for that season. As shown in the figure, we obtained a large number of spectra during Season 2022 (1106 over 39 nights), which enabled us to perform a frequency analysis over a broad range with high confidence. The lower limit 0.005 d-1 corresponds to $1/T$, where T is the observing period (194.7 days), and the upper limit is 1.2 d-1.
As with other analyses, we started with the simple GLS periodogram without pre-whitening. The corresponding power spectrum is presented in Fig. 11(a). We determined that all the significant frequencies lie below 1.2 d-1. As in the previous analysis, we determined the first seven frequencies (see Table 6). A period of $\sim 10$ days is also observed, taking into account the uncertainties in this season. The most powerful period of the spectroscopic data is $\sim 16.2$ days. It should be noted that a general view of the periodogram is different from the power spectrum of the TESS photometry (Fig. 5).
Table 6: Periods detected in Season 2022.111111Results of the analysis using three different methods, GLS, GLSp, and WWZ, arranged by decreasing period in each column. The number in brackets [N] denotes the rank of the period by detected power.
@@
To derive the periods and build the phase-folded diagrams, we performed the GLSp; the resulting periods are listed in Table 6. We note that $F_{\mathrm{5}}$ $\sim$ $F_{\mathrm{2}}$ + $F_{\mathrm{3}}$. In Fig. 11(b), we present a phase plot with the first most powerful period we found. The curve exhibits a pronounced sawtooth shape.
@@
Figure 12: Result of the WWZ analysis for Season 2022. The wavelet power is indicated by the colour (see colour bar at left). The black solid line in the right panel shows the time-averaged wavelet power.
Lastly, we obtained a very interesting result using the WWZ analysis (Fig. 12 and Table 6). The frequency range was taken from 0.01 to 0.25 d-1. The lowest and strongest frequency on the scalogram corresponds to the period $\sim 16.4$ days. According to the scalogram, this period was very pronounced and existed throughout the entire observation season and slightly increases by the end of the season. The period of $\sim 11$ days also exists in the scalogram of Season 2022, and here the value is equal to 10 $\pm$ 1 days, although this period does not exist all the time and increases towards the end, according to the scalogram.